# Structure of the Cayley-Dickson algebras

### v1

@@ Abstract
Viewing the Cayley-Dickson process as a graded construction provides a rigorous definition of associativity consisting of three types and the non-associative parts dividing into four types. These simplify the Moufang loop identities and Mal’cev’s identity, which identifies the non-associative Lie algebra structure. This structure is distinct for the first four power-associative algebras and has the same pattern for at least the next three. The structure is identified with 3-cycles and modes that reduce sets of 84 zero divisors to just 7, in most cases, such as for the sedenions, and identifies the subalgebas of the power associative algebras that provide zero divisors thus defining the structure of Cayley-Dickson algebras. The zero divisor and octonion cardinality at all levels is derived. Split algebras are proved to be isomorphic for any power-associative algebra and share the same non-associative structure with different zero divisors. Due to the inclusion of new power-associative algebras into the hypercomplex algebras the terminology ultracomplex numbers is suggested.
Keywords. Octonions, sedenions, Cayley-Dickson algebras, split-sedenions, zero divisors, graded algebras, Moufang loops, Malcev algebra, ultracomplex numbers.
@@ 1 Introduction
The Cayley-Dickson construction generates a series of algebras starting with the only four normed division algebras, the reals, complex numbers, quaternions and octonions, denoted ${\mathbb{R}}$, ${\mathbb{C}}$, ${\mathbb{H}}$ and ${\mathbb{O}}$, respectively [1, 2]. After this is an infinite series of power-associative algebras whose properties are analysed herein using a graded notation. These algebras are generally not considered to be graded because pure basis elements have the same properties so permutations are presumed to be isomorphic to each other. Also each level of the construction involves the product of the two Cayley-Dickson algebras at the previous level, which demonstrates that each algebra contains the previous algebra as a subalgebra. This statement and many other properties become obvious once a graded notation is adopted because the Cayley-Dickson process is naturally suited to be graded. The first power-associative algebra is the sedenions, which are not a division algebra due to the existence of zero divisors. The Cayley-Dickson construction includes the sedenions at all higher levels but the associativity properties are different providing structure that is found to be unique for the first four power-associative algebras. Thereafter it is proved that the structure can not become more complicated, just increasing in size so all remaining power-associative algebras have the same structure.
In ungraded form the Cayley-Dickson construction is [3],
@@
where the conjugate is $(a,b)^{*}=(a^{*},-b)$ and $\epsilon=1$ for the normed composition algebras analysed by Hurwitz [4]. The internal products, $ac$ for instance, are then expanded at the next level and the process repeated generating a polynomial binomial expansion pyramid easily represented by a computer program. This process stops at the reals, ${\mathbb{R}}$. Each new level is generally shown with a new generator operating on the previous level. So quaternions extend $i$ with $j$ providing a new element $k=ij$, which makes the notation ungraded. For octonions, another basis element is introduced, generally $l$, which extends the quaternions in a mixed graded and ungraded notation [4]. A consistent graded notation is obviously needed and the benefits allow the series to be analysed to arbitrary levels, which uncovers different types of associativity that are only exposed in the power-associative algebras.
@@
Deriving an ordering using the graded notation finds that power-associative algebras have three types of associativity, which leads to four types of non-associativity. Octonions only have one symmetric type of non-associativity that makes the three Moufang loop identities identical. Sedenions introduce two more asymmetric non-associative types, which allows for zero divisors and the second power-associative algebra, called the trigintaduonions, includes the final non-associative type which adds more complicated zero divisors. Analysis of the new non-associative types uncovers cyclic forms and modes that categorise zero divisors and applied to the 84 known sedenion zero divisors [5], reduces them to seven distinct triples of basis elements with three cycles and four modes each. The trigintaduonions have 1,260 zero-divisors, which are derived from 147 distinct triples. Zero divisor cardinality is derived as multiples of 84 and copies of ${\mathbb{O}}$ subalgebras are provided. The zero divisors belong to other algebras related to octonions but containing zero divisors hence introducing power associativity into the Cayley-Dickson series of algebras. Analysing these new pseudo-octonion algebras derives the structure of the infinite series. It is found that the pseudo-octonion algebras increase at a fixed rate and the octonions are like remainders. Octonions are the first step in the series existing as a single remainder.
@@
The next section starts by building the Cayley-Dickson construction as two equivalent pyramids using graded and binary notations. Swapping between the two allows the usual composition properties to be proved for all Cayley-Dickson algebras showing that the graded notation is indeed equivalent to the usual doubling of algebras [6]. The structure section derives the four non-associativity types and the eight 3-cycles that present the structure that simplifies the Moufang identities and Malcev algebras. The Malcev algebra defines non-associative Lie algebras and this is used to count subalgebras isomorphic to octonions and some of the power-associative pseudo-algebras. The structure of Cayley-Dickson algebras is derived from the relationship between these subalgebras and octonions. The fourth section defines zero divisors and modes and shows the relationship to the eight 3-cycles. It proceeds to document the seven and 147 prime zero divisors of the first two power-associative algebras and applies the Cycles and Modes theorems to uncover the rest. The final section shows each power-associative algebra has a single split version with the same non-associative structure but different zero divisors. These are provided for the split sedenions.
@@ 2 Graded Cayley-Dickson Construction
The basis of the Cayley-Dickson algebra at level $n$, called the ungraded dimension, has $2^{n}$ graded elements consisting of grades of size $\binom{n}{k}$ called a $k$-grade. Basis elements are represented as $o_{\alpha}$ where $\alpha$ is a set of digits from ${\mathrm{A}(n)}=\{\emptyset,(1),(12),\ldots,(12\cdots n)\}$, generated by the recurrence relation ${\mathrm{A}(n)}={\mathrm{A}(n-1)}+{\mathrm{A}(n-1)}(n)$ where set multiplication by $(n)$ extends each member of the set by $n$, ${\mathrm{A}(n-1)}(n)=\{(n),(1n),(12n),\ldots,(12\cdots(n-1)n)\}$. This operation is called expansion but multiplying by an existing member will contract some members. Using the notation of Porteous [7] whereby ${\mathbb{N}}_{1}^{n}$ is the set of natural numbers from 1 up to and including $n$, then $\{(12\cdots n)\}(k)=\{(12\cdots(k-1)(k+1)\cdots n)\}$ for $k\in{\mathbb{N}}_{3}^{n-1}$. Note that ${\mathrm{A}(n)}$ includes the empty set, $\emptyset$, or no digit case, to denote the unit graded element, 1, which is the only graded element at level zero and the unit notation is preferred to the use of the empty set, $o=1$. The set brackets are dropped in the graded element, $(\alpha)\rightarrow o_{\alpha}$.
@@
A pure element excludes the unit basis element, 1. A blade is a pure, non-zero, scaled basis element and a $k$-blade is a member of the $k$-grade. Note that blades may have positive or negative squares and for the remainder this assumed to be ±$\pm 1$, for blades. A blade is positive or negative for scale +1 or -1, respectively.
@@
The pyramid element ${a}$, at level $n$, specifies the $n^{\rm th}$ row of a binary pyramid with coefficients $a_{\beta}$, $\beta\in{\mathrm{B}(n)}$, where ${\mathrm{B}(n)}={\mathrm{B}(n-1)}(0)+{\mathrm{B}(n-1)}(1)$ so that ${\mathrm{B}(n)}$ is all binary numbers with $n$ digits from ${\mathbb{N}}_{0}^{1}$ and set multiplication extends the binary numbers in ${\mathrm{B}(n-1)}$. Hence ${\mathrm{B}(n)}$ covers all $2^{n}$ combinations of $n$ digits in ${\mathrm{A}(n)}$ and this is an equivalence relationship to graded elements with coefficients $a_{\beta}=0$ or ±$\pm 1$ because appending a $1$ to $\beta$ at the $n$th position is the same as appending $n$ to the graded recurrence relation. Hence, at level 1, ${a}_{0}=(\pm 1,0)=\pm 1$ and ${a}_{1}=(0,\pm 1)=\pm o_{1}$. To level three this is
@@
±$\begin{split}\pm o_{1}&=(((0,\pm 1),(0,0)),((0,0),(0,0)))\text{ and}\\ \pm o_{123}&=(((0,0),(0,0)),((0,0),(0,\pm 1)))\text{.}\end{split}$
@@
The Cayley-Dickson product at level $i$ for two pure pyramid elements, ${a}_{\alpha}{b}_{\beta}$, with $\alpha,\beta\in{\mathrm{B}(n)}$, with $\alpha_{i}$ being all indices to the left of position $i$ and $\alpha_{{\overline{i}}}$ being all indices to the right of position $i$, is
$\begin{split}(a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}},a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}})(b_{\beta_{i}{0}\beta_{\hskip 1.0pt{\overline{i}}}},b_{\beta_{i}{1}\beta_{\hskip 1.0pt{\overline{i}}}})&=\\ \Big{(}a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}}b_{\beta_{i}{0}\beta_{\hskip 1.0pt{\overline{i}}}}&-\epsilon_{i}b^{*}_{\beta_{i}{1}\beta_{\hskip 1.0pt{\overline{i}}}}a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}},\>b_{\beta_{i}{1}\beta_{\hskip 1.0pt{\overline{i}}}}a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}}+a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}b^{*}_{\beta_{i}{0}\beta_{\hskip 1.0pt{\overline{i}}}}\Big{)}\text{,}\end{split}$ (3)
where $\epsilon_{i}=\pm 1$ for $i\in{\mathbb{N}}_{1}^{n}$ and the conjugate operator at level $i$ is $a^{*}_{\alpha}0=(a^{*}_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}},-a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}})$ and the conjugate operator on the right hand side acts at level $i-1$ and has no effect at level 0. Equation (3) corresponds to a section of the pyramid one pair deep at level $i$. The terms on the right hand side of (3) are called the parallel and cross terms, respectively. The $\alpha_{{\overline{i}}}$ indices are not shown if empty.
@@
For $n\geq 2$, showing just the first pair in each half, with $\alpha\in{\mathrm{B}(n)}$ and $\beta=00\cdots 0$ of length $n-2$, then
@@
$\begin{split}&(a_{0},a_{1})(b_{0},b_{1})\\ &=\big{(}(a_{00},a_{10}),(a_{01},a_{11}))((b_{00},b_{10}),(b_{01},b_{11})\big{)}\\ &=\big{(}(a_{00},a_{10})(b_{00},b_{10})-\epsilon_{2}(b_{01},b_{11})^{*}(a_{01},a_{11}),\\ &\enspace\enspace\enspace\enspace(b_{01},b_{11})(a_{00},a_{10})+(a_{01},a_{11})(b_{00},b_{10})^{*}\big{)}\\ &=\big{(}(a_{00}b_{00}-\epsilon_{1}b_{10}a_{10},b_{10}a_{00}+a_{10}b_{10})-\epsilon_{2}(b_{01}a_{01}+\epsilon_{1}a_{11}b_{11},b_{01}a_{11}-b_{11}a_{01}),\\ &\enspace\enspace\enspace\enspace\big{(}b_{01}a_{00}-\epsilon_{1}a_{10}b_{11},a_{10}b_{01}+b_{11}a_{00})+(a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},-b_{10}a_{01}+a_{11}b_{00})\big{)}\\ &=\big{(}(a_{00}b_{00}-\epsilon_{1}b_{10}a_{10}-\epsilon_{2}b_{01}a_{01}-\epsilon_{1}\epsilon_{2}a_{11}b_{11},b_{10}a_{00}+a_{10}b_{00}-\epsilon_{2}b_{01}a_{11}+\epsilon_{2}b_{11}a_{01}),\\ &\enspace\enspace\enspace\enspace\big{(}b_{01}a_{00}-\epsilon_{1}a_{10}b_{11}+a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},a_{10}b_{01}+b_{11}a_{00}-b_{10}a_{01}+a_{11}b_{00})\big{)}\text{.}\end{split}$ (6)
@@
$\begin{split}\sum_{\beta}(a_{\beta 0},a_{\beta 1})(b_{\beta 0},b_{\beta 1})&=((a_{00}b_{00}-\epsilon_{1}b_{10}a_{10}-\epsilon_{2}b_{01}a_{01}-\epsilon_{1}\epsilon_{2}a_{11}b_{11},0),\\[-12.91663pt] &\enspace\enspace\enspace\enspace(b_{01}a_{00}-\epsilon_{1}a_{10}b_{11}+a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},0))\text{.}\end{split}$
@@
$a_{\alpha}b_{\alpha}=a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}b_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}=(-\epsilon_{i}b^{*}_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}},0)$
@@
[…]
The first case is positive because the original conjugation negated the $o_{i}$ element and a new conjugate for the $i-1$ level was introduced whereas for $b^{*}_{\alpha_{i-1}{0}\alpha_{\hskip 1.0pt{\overline{i-1}}}}$ the conjugation is passed to the next level without effect. Continuing to the following levels has no effect on the original sign and together will introduce $\epsilon_{j}$ from $a^{2}_{{\alpha_{j}{1}\alpha_{\hskip 1.0pt{\overline{j}}}}}\forall j\in i_{1},i_{2},\ldots,i_{k-1}$. This means $o_{i_{1}i_{2}\ldots i_{k}}^{2}=-\epsilon_{i_{1}}\epsilon_{i_{2}}\cdots\epsilon_{i_{k}}$ which is assumed true and by induction $a^{2}_{\alpha 1}$ is considered for $\alpha$ having ones in positions $i_{1}i_{2}\ldots i_{k}$ and zeros elsewhere. Then
@@
The Cayley-Dickson construction to any level, $n$, generates an algebra with pure elements, $x$, $y$ satisfying $xy=-yx$ if $x\neq y$ and, if $\epsilon_{i}=1$, $\forall i\in{\mathbb{N}}_{1}^{n}$ and $x=y$, then $xy=-1$. The unit, 1, has no effect under multiplication.
@@
$(a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}},a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}})(b_{\beta},0)$ $=\left(a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}}b_{\beta},a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}b^{*}_{\beta}\right)$
$=\big{(}(a_{\alpha_{i-1}{0}\alpha_{\hskip 1.0pt{\overline{i-1}}}},0)(b_{\beta},0),(a_{\alpha_{i-1}{1}\alpha_{\hskip 1.0pt{\overline{i-1}}}},0)(b_{\beta},0)\big{)}\text{.}$

### v2

@@ Abstract
Viewing the Cayley-Dickson process as a graded construction provides a rigorous definition of associativity consisting of three types and the non-associative parts dividing into four types. These simplify the Moufang loop identities and Mal’cev’s identity, which identifies the non-associative Lie algebra structure. Many authors have assumed the Moufang loop identities are identical but this is not the case for sedenions. Analysing the non-associativity structure uncovers 3-cycles that distinguish between the Moufang identities and are used to identify three power-associative subalgebras of sedenions and higher level Cayley-Dickson algebras.
Power-associativity introduces zero divisors into Cayley-Dickson algebras in a systematic way and it is convenient to replace the terminology hypercomplex numbers with ultracomplex numbers for the power-associative algebras. The non-associative types simplify the Moufang identities and show that zero divisors in these algebras occur in multiples of 84. Modes are uncovered that reduce these down to factors of seven base zero divisor pairs. It is shown that this is due to the the power-associative subalgebras being embedded into ultracomplex numbers in multiples of seven. The graded notation allows the eight octonion and seven power-associative subalgebras of sedenions to be uniquely derived, up to representation.
Split Cayley-Dickson algebras are proved to be isomorphic to each other for any ultracomplex algebra and these share the same non-associative structure as the non-split case but with different zero divisors. These are provided for the split sedenions.
Keywords. Octonions, sedenions, Cayley-Dickson algebras, split sedenions, zero divisors, graded algebras, Moufang loops, Malcev algebra, ultracomplex numbers.
@@ 1 Introduction
The standard approach to the Cayley-Dickson construction [5] at step $n$ is to define a multiplication operation for $A_{n}\cong A_{n-1}\oplus A_{n-1}$ by
@@
where the conjugate is $(a,b)^{*}=(a^{*},-b)$ and $\epsilon=1$ for the normed composition algebras analysed by Hurwitz [6]. The products $ac$, $d^{*}b$, $da$ and $bc^{*}$ are defined according to the construction for $A_{n-1}\cong A_{n-2}\oplus A_{n-2}$, which in turn is eventually determined by elements in $A_{0}={\mathbb{R}}$.
Traditionally, the multiplication operator is introduced using anticommuting imaginaries $i,j,l$ for the complex numbers, quaternions and octonions, respectively, as
$A_{1}$ $={\mathbb{R}}\oplus i{\mathbb{R}}={\mathbb{C}}\text{,}$
$A_{2}$ $={\mathbb{C}}\oplus{\mathbb{C}}j={\mathbb{H}}\text{ and}$
$A_{3}$ $={\mathbb{H}}\oplus l{\mathbb{H}}={\mathbb{O}}\text{.}$ (1)
This paper introduces a graded decomposition of anticommuting elements, $o_{n}$, at each level, $n$, as
$A_{n}=A_{n-1}\oplus A_{n-1}o_{n}$
so that $o_{1}=i$, $o_{2}=j$ and $o_{3}=-l$. The ungraded notation of using $k=ij$ is then represented in graded form as $ij=o_{12}$. All elements of $A_{n}$ are generated using the binomial expansion, which provides an ordering of these elements. This is not important for the normed division algebras and, for example, the exception Lie algebra G2, which are the automorphisms for octonions. Swapping any pair of the seven elements of octonions generates another representation of octonions. The ordering enables power-associativity but only comes into play for $A_{n}$ with $n\geq 4$ where $A_{4}={\mathbb{S}}$, the sedenions.
@@
Deriving an ordering using the graded notation finds that power-associative algebras have three types of associativity, which leads to four types of non-associativity. Octonions only have one symmetric type of non-associativity that makes the three Moufang loop identities identical. Sedenions introduce two more non-associative types, which separates two of the Moufang identities and allows for zero divisors. The trigintaduonions, $A_{5}$, include the final non-associative type, identified by the remaining Moufang identity giving unique results, which adds more complicated zero divisors. Using the new non-associative types provides a simpler categorisation than the Moufang identities. Analysis of the new non-associative types uncovers cyclic forms and modes that categorise zero divisors and applied to the 84 known sedenion zero divisors [7], reduces them to seven distinct triples of basis elements with three cycles and four modes each. The trigintaduonions have 1,260 zero-divisors, which are derived from 147 distinct triples. The cyclic forms identify the octonion and three power-associative sub-algebras of $A_{5}$, called pseudo-octonion algebras in [10]. Each of these algebras contain 12 zero divisors and since they are embedded in $A_{n}$, $n\geq 4$, in multiples of seven, then it is found that zero divisors in $A_{n}$, $n>3$ occur in multiples of 84. The cardinality formula for zero divisors for all $A_{n}$ is derived.
Using loop analysis, Cawagas [7] has previously identified the first power-associative subalgebra by identifying that it did not satisfy one of the three Moufang identities, unlike octonions. He called this quasi-octonions and provided an enumeration in [9]. This paper takes a different approach to loop analysis to prove that most of the sedenion triple products satisfy the three Moufang identities in different ways. Previous work [10] has shown that the quasi-octonions are part of a series of six algebras, called pseudo-octonions to allow for the designation $\mathbb{P}_{k}$}, $k\in\{4,8,10,12,14,16\}$ where $k$ is the number of distinct non-associative triple products. These algebras where shown to break the symmetry of octonions in a way consistent with exceptional Lie algebra G2. Cawagas found seven copies of $\mathbb{P}_{4}$} embedded in sedenions as well as eighth copies of octonions. Another two of these power-associative algebras are introduced in the next two levels of Cayley-Dickson algebras, $A_{5}$ and $A_{6}$.
The octonions have been found to apply to areas in particle physics [11, 12] and extended to sedenions more recently [13, 14, 15, 16]. It may not be necessary to go to higher dimensions to achieve new representations in particle physics. All Cayley-Dickson algebras of dimension higher than octonions consist of multiple copies of octonions and various pseudo-octonions so it is possible that just octonions and the pseudo-octonion algebras are all that is necessary.
@@
Section 2 starts by building the Cayley-Dickson construction as two equivalent pyramids using graded and binary notations. Swapping between the two allows the usual composition properties to be proved for all Cayley-Dickson algebras showing that the graded notation is indeed equivalent to the usual doubling of algebras [8]. Section 3 derives the four non-associativity types that simplify the Moufang identities and Malcev algebras and derives the eight non-associative 3-cycles and 3-cycle structure in the Cycle theorem that identifies the subalgebras of $A_{n}$ for $n\geq 3$. The 15 octonion and $\mathbb{P}_{4}$} representations embedded in sedenions are identified using the graded notation.
Section 4 defines zero divisors and modes in the Mode theorem and shows the relationship to the eight 3-cycles. It proceeds to document the seven and 147 prime zero divisors of $A_{4}$ and $A_{5}$, respectively, and applies the Cycles and Modes theorems to uncover the rest. Section 5 shows each Cayley-Dickson algebra has a single split version with the same non-associative structure but different zero divisors. These are provided for the split sedenions.
@@ 2 Graded Cayley-Dickson Construction
The basis of the Cayley-Dickson algebra at level $n$, called the ungraded dimension, has $2^{n}$ graded elements consisting of grades of size $\binom{n}{k}$ called a $k$-grade. Basis elements are represented as $o_{\alpha}$ where $\alpha$ is a set of digits from ${\mathrm{A}(n)}=\{\emptyset,(1),(12),\ldots,(12\cdots n)\}$, generated by the recurrence relation ${\mathrm{A}(n)}={\mathrm{A}(n-1)}+{\mathrm{A}(n-1)}(n)$, where set multiplication by $(n)$ extends each member of the set by $n$, ${\mathrm{A}(n-1)}(n)=\{(n),(1n),(12n),\ldots,(12\cdots(n-1)n)\}$. This operation is called expansion but multiplying by an existing member will contract some members. Using the notation of Porteous [17] whereby ${\mathbb{N}}_{1}^{n}$ is the set of natural numbers from 1 up to and including $n$, then $\{(12\cdots n)\}(k)=\{(12\cdots(k-1)(k+1)\cdots n)\}$ for $k\in{\mathbb{N}}_{3}^{n-1}$. Note that $\mathrm{A}(n)$} includes the empty set, $\emptyset$, or no digit case, to denote the unit graded element, 1, which is the only graded element at level zero and the unit notation is preferred to the use of the empty set, $o=1$. The set brackets are dropped in the graded element, $(\alpha)\rightarrow o_{\alpha}$.
@@
A pure element excludes the unit basis element, 1. A blade is a pure, non-zero, scaled basis element and a $k$-blade is a member of the $k$-grade. Note that blades may have positive or negative squares and for the remainder this assumed to be $\pm 1$, for blades. A blade is positive or negative for scale +1 or -1, respectively.
@@
The pyramid element a{a}, at level $n$, specifies the $n^{\rm th}$ row of a binary pyramid with coefficients $a_{\beta}$, $\beta\in{\mathrm{B}(n)}$, where ${\mathrm{B}(n)}={\mathrm{B}(n-1)}(0)+{\mathrm{B}(n-1)}(1)$ so that $\mathrm{B}(n)$} is all binary numbers with $n$ digits from ${\mathbb{N}}_{0}^{1}$ and set multiplication extends the binary numbers in $\mathrm{B}(n-1)$}. Hence $\mathrm{B}(n)$} covers all $2^{n}$ combinations of $n$ digits in $\mathrm{A}(n)$} and this is an equivalence relationship to graded elements with coefficients $a_{\beta}=0$ or $\pm 1$ because appending a $1$ to $\beta$ at the $n$th position is the same as appending $n$ to the graded recurrence relation. Hence, at level 1, ${a}_{0}=(\pm 1,0)=\pm 1$ and ${a}_{1}=(0,\pm 1)=\pm o_{1}$. To level three this is
@@
$\begin{split}\pm o_{1}&=(((0,\pm 1),(0,0)),((0,0),(0,0)))\text{ and}\\ \pm o_{123}&=(((0,0),(0,0)),((0,0),(0,\pm 1)))\text{.}\end{split}$
@@
The Cayley-Dickson product at level $i$ for two pure pyramid elements, ${a}_{\alpha}{b}_{\beta}$, with $\alpha,\beta\in{\mathrm{B}(n)}$, with $\alpha_{i}$ being all indices to the left of position $i$ and $\overline{\alpha_{i}}$} being all indices to the right of position $i$, is
$\begin{split}(a_{\alpha_{i}{0}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt},a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt})(b_{\beta_{i}{0}{\overline{\beta_{i}}}\hskip 1.0pt},b_{\beta_{i}{1}{\overline{\beta_{i}}}\hskip 1.0pt})&=\\ \Big(a_{\alpha_{i}{0}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}b_{\beta_{i}{0}{\overline{\beta_{i}}}\hskip 1.0pt}&-\epsilon_{i}b^{*}_{\beta_{i}{1}{\overline{\beta_{i}}}\hskip 1.0pt}a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt},\>b_{\beta_{i}{1}{\overline{\beta_{i}}}\hskip 1.0pt}a_{\alpha_{i}{0}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}+a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}b^{*}_{\beta_{i}{0}{\overline{\beta_{i}}}\hskip 1.0pt}\Big)\text{,}\end{split}$ (4)
where $\epsilon_{i}=\pm 1$ for $i\in{\mathbb{N}}_{1}^{n}$ and the conjugate operator at level $i$ is $a^{*}_{\alpha}0=(a^{*}_{\alpha_{i}{0}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt},-a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt})$ and the conjugate operator on the right hand side acts at level i1i-1 and has no effect at level 0. Equation (4) corresponds to a section of the pyramid one pair deep at level $i$. The terms on the right hand side of (4) are called the parallel and cross terms, respectively. The $\overline{\alpha_{i}}$} indices are not shown if empty.
@@
For $n\geq 2$, showing just the first pair in each half, with $\alpha\in{\mathrm{B}(n)}$ and $\beta=00\cdots 0$ of length n2n-2, then
@@
$\begin{split}&(a_{0},a_{1})(b_{0},b_{1})\\ &=\big((a_{00},a_{10}),(a_{01},a_{11}))((b_{00},b_{10}),(b_{01},b_{11})\big)\\ &=\big((a_{00},a_{10})(b_{00},b_{10})-\epsilon_{2}(b_{01},b_{11})^{*}(a_{01},a_{11}),\\ &\kern 5.0pt\kern 5.0pt\kern 5.0pt\kern 5.0pt(b_{01},b_{11})(a_{00},a_{10})+(a_{01},a_{11})(b_{00},b_{10})^{*}\big)\\ &=\big((a_{00}b_{00}-\epsilon_{1}b_{10}a_{10},b_{10}a_{00}+a_{10}b_{10})-\epsilon_{2}(b_{01}a_{01}+\epsilon_{1}a_{11}b_{11},b_{01}a_{11}-b_{11}a_{01}),\\ &\kern 5.0pt\kern 5.0pt\kern 5.0pt\kern 5.0pt\big(b_{01}a_{00}-\epsilon_{1}a_{10}b_{11},a_{10}b_{01}+b_{11}a_{00})+(a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},-b_{10}a_{01}+a_{11}b_{00})\big)\\ &=\big((a_{00}b_{00}-\epsilon_{1}b_{10}a_{10}-\epsilon_{2}b_{01}a_{01}-\epsilon_{1}\epsilon_{2}a_{11}b_{11},b_{10}a_{00}+a_{10}b_{00}-\epsilon_{2}b_{01}a_{11}+\epsilon_{2}b_{11}a_{01}),\\ &\kern 5.0pt\kern 5.0pt\kern 5.0pt\kern 5.0pt\big(b_{01}a_{00}-\epsilon_{1}a_{10}b_{11}+a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},a_{10}b_{01}+b_{11}a_{00}-b_{10}a_{01}+a_{11}b_{00})\big)\text{.}\end{split}$ (7)
@@
$\begin{split}\sum_{\beta}(a_{\beta 0},a_{\beta 1})(b_{\beta 0},b_{\beta 1})&=((a_{00}b_{00}-\epsilon_{1}b_{10}a_{10}-\epsilon_{2}b_{01}a_{01}-\epsilon_{1}\epsilon_{2}a_{11}b_{11},0),\\[-12.91663pt] &\kern 5.0pt\kern 5.0pt\kern 5.0pt\kern 5.0pt(b_{01}a_{00}-\epsilon_{1}a_{10}b_{11}+a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},0))\text{.}\end{split}$
@@
$a_{\alpha}b_{\alpha}=a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}b_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}=(-\epsilon_{i}b^{*}_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt},0)$
@@
$b^{*}_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}=\begin{cases}(\epsilon_{i-1}b^{*}_{\alpha_{i-1}{1}\hskip 1.0pt{\overline{\alpha_{i-1}}}\hskip 1.0pt}a_{\alpha_{i-1}{1}\hskip 1.0pt{\overline{\alpha_{i-1}}}\hskip 1.0pt},0)&\text{if $i-1\in i_{1},i_{2},\ldots,i_{k}$,}\\ (b^{*}_{\alpha_{i-1}{0}\hskip 1.0pt{\overline{\alpha_{i-1}}}\hskip 1.0pt}a_{\alpha_{i-1}{0}\hskip 1.0pt{\overline{\alpha_{i-1}}}\hskip 1.0pt},0)&\text{otherwise}\text{.}\end{cases}$
The first case is positive because the original conjugation negated the $o_{i}$ element and a new conjugate for the i1i-1 level was introduced whereas for $b^{*}_{\alpha_{i-1}{0}\hskip 1.0pt{\overline{\alpha_{i-1}}}\hskip 1.0pt}$ the conjugation is passed to the next level without effect. Continuing to the following levels has no effect on the original sign and together will introduce $\epsilon_{j}$ from $a^{2}_{{\alpha_{j}{1}\hskip 1.0pt{\overline{\alpha_{j}}}\hskip 1.0pt}}\forall j\in i_{1},i_{2},\ldots,i_{k-1}$. This means $o_{i_{1}i_{2}\ldots i_{k}}^{2}=-\epsilon_{i_{1}}\epsilon_{i_{2}}\cdots\epsilon_{i_{k}}$ which is assumed true and by induction $a^{2}_{\alpha 1}$ is considered for $\alpha$ having ones in positions $i_{1}i_{2}\ldots i_{k}$ and zeros elsewhere. Then
@@
The Cayley-Dickson construction to any level, $n$, generates an algebra with pure elements, $x$, $y$ satisfying xy=yxxy=-yx if $x\neq y$ and, if $\epsilon_{i}=1$, $\forall i\in{\mathbb{N}}_{1}^{n}$ and $x=y$, then xy=1xy=-1. The unit, 1, has no effect under multiplication.
@@
$(a_{\alpha_{i}{0}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt},a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt})(b_{\beta},0)$ $=\left(a_{\alpha_{i}{0}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}b_{\beta},a_{\alpha_{i}{1}\hskip 1.0pt{\overline{\alpha_{i}}}\hskip 1.0pt}b^{*}_{\beta}\right)$
$=\big((a_{\alpha_{i-1}{0}\hskip 1.0pt{\overline{\alpha_{i-1}}}\hskip 1.0pt},0)(b_{\beta},0),(a_{\alpha_{i-1}{1}\hskip 1.0pt{\overline{\alpha_{i-1}}}\hskip 1.0pt},0)(b_{\beta},0)\big)\text{.}$